OpenStudy (johnnydicamillo):

f'(x) where f(x) = ln((ex^2)/(x-5)^7)

OpenStudy (johnnydicamillo):

\[f(x) = \ln(\frac{ ex^2 }{ (x-5)^7 })\]

OpenStudy (freckles):

Use properties of log first

OpenStudy (freckles):

That will make the differentiation process tons easier

OpenStudy (johnnydicamillo):

so\[\frac{ lnex^2 }{ \ln (x-5)^7 }\]

OpenStudy (freckles):

That isn't a property of log

OpenStudy (johnnydicamillo):

Can I have a hint?

OpenStudy (johnnydicamillo):

\[ex^2 \ln (x-5)^7\]

OpenStudy (freckles):

My hint is that you review the properties of log

OpenStudy (freckles):

ln(a/b) =ln(a) - ln(b)

ganeshie8 (ganeshie8):

\[\large \rm \ln\left(\dfrac{\spadesuit }{\heartsuit }\right) = \ln\left(\spadesuit\right)-\ln\left(\heartsuit\right)\]

ganeshie8 (ganeshie8):

\[\large \rm \ln\left(\spadesuit \heartsuit \right) = \ln\left(\spadesuit\right)+\ln\left(\heartsuit\right)\]

OpenStudy (freckles):

http://www.purplemath.com/modules/logrules.htm

OpenStudy (johnnydicamillo):

okay so ln (ex^2)- ln (x-5)^7

ganeshie8 (ganeshie8):

you can work this without logs if u dont like them, bit it would be very messy using quitient and product rules

OpenStudy (freckles):

You can even apply the power rule for log

OpenStudy (jhannybean):

lmao, spades and hearts huh?

OpenStudy (johnnydicamillo):

2ln ex - 7 ln (x-5)

OpenStudy (johnnydicamillo):

How far off am I

OpenStudy (freckles):

Good job

OpenStudy (freckles):

Wait was that 2 on the ex or the x

OpenStudy (johnnydicamillo):

its \[ex^2\]

OpenStudy (freckles):

So the second term you have is right but the first term is a bit off

ganeshie8 (ganeshie8):

lol i got those from satellite i think... hey \(\large \ln (ex^2) = \ln(e) + \ln(x^2)\)

OpenStudy (jhannybean):

if it's \(ex^2\) then e is a function and so is \(x^2\), therefore product rule is applied.

OpenStudy (jhannybean):

Oh log rule of the numerator alone. It's like inception.

OpenStudy (johnnydicamillo):

haha so ln(e) + ln (x^2) - 7 ln (x -5)

OpenStudy (jhannybean):

Seems better. Does everyone else agree?

OpenStudy (johnnydicamillo):

\[\frac{ 1 }{ e} + \frac{ 1 }{ x^2 } - 7 \frac{ 1 }{ x-5 }\]

OpenStudy (johnnydicamillo):

do I need chain rule?

OpenStudy (freckles):

Yea we can do one more thing but I think this should be easy as is already to differentiate

OpenStudy (jhannybean):

you're partially right there

OpenStudy (jhannybean):

With log functions, you have a function inside of a function, you most always use chain rule

OpenStudy (freckles):

ln(e) is a constant

OpenStudy (jhannybean):

There's exceptions sometimes.

OpenStudy (johnnydicamillo):

so if ln(e) is a constant that is = 0

OpenStudy (jhannybean):

Remember, \(\ln (e) = \log_{e} e = 1\)

OpenStudy (freckles):

And you need to look at your second term also

OpenStudy (johnnydicamillo):

\[\frac{ 1 }{ x^2} * 2x - 7 \frac{ 1 }{ x - 5 } * 1\]

OpenStudy (jhannybean):

\[\frac{ 1 }{ e} + \frac{ 1 }{ x^2 } - 7 \frac{ 1 }{ x-5 }\] first part: constant - deirvative = 0 second part: you still got an x^2 - take its derivative, 3rd part - x-5 is sill a function, and the derivative of x =1, so what are you going to do now

OpenStudy (johnnydicamillo):

Is it what I posted above?

OpenStudy (jhannybean):

\[\frac{ 1 }{ e} + \frac{ 1 }{ x^2 } - 7 \frac{ 1 }{ x-5 }\]\[0 + \frac{1}{x^2} \cdot 2x -\frac{7}{x-5}\cdot 1\]\[\frac{2x}{x^2} - \frac{7}{x-5}\]\[2x^{1-2} -\frac{7}{x-5}\]\[2^{-1} -\frac{7}{x-5}\]\[\boxed{\large \frac{2}{x} -\frac{7}{x-5}}\]There you have it

OpenStudy (jhannybean):

You were almost there. You have any questions about it?

OpenStudy (jhannybean):

not sure if you have to totally simplify it either with the common denominator and stuff, that would make it look pretty but it'd hurt.

OpenStudy (johnnydicamillo):

I did \[2x/(x^2)-7/(x-5)\]

OpenStudy (jhannybean):

Well, you don't want to leave it like that because the x's inn the first fraction can indeed simplify further. What I meant was, once you have your boxxed answer, you can actually simplify it further by obtaining the LCD of the two functions, and this and that.

OpenStudy (johnnydicamillo):

Okay, well thanks for your assistance.

OpenStudy (jhannybean):

No problem. IF you are interested in simplifying it further, it would be like this:\[\frac{2}{x} -\frac{7}{x-5}\]\[\frac{2x-10}{x^2 -5x} -\frac{7x}{x^2-5x}\]\[\boxed{\large \frac{-9x-10}{x^2-5x}}\]